To:  Roger Sawyer, His Nibs
From:  George McNulty

Professor and Humble Servants

Subject: Subjection
Date:  32 March 1999
To: Roger Sawyer, His Nibs
From: George McNulty

Professor and Humble Servants

Subject: Subjection
Date: 32 March 1999
To:  Roger Sawyer, His Nibs
From:  George McNulty

Professor and Humble Servants

Subject:  Subjection
Date:  32 March 1999

Now that the header information is out of the way, let’s get on with the show. And this is a second paragraph. Note that a real memo would have extra space between paragraphs in addition to not indenting.

Name  Salary
George McNulty $255,492,100.00
Doug Meade  $ 49.47
Name  Salary.
George McNulty $255,492,100.00
Doug Meade  $ 49.47
Name  Salary
George McNulty $255,492,100.00
Doug Meade  $ 49.47
Name  Salary
George McNulty $255,492,100.00
Doug Meade  $ 49.47
Name  Salary
George McNulty $255,492,100Doug00
Doug Meade  $ 49Doug47
Name  Salary
George McNulty $255,492,100.00
Doug Meade  $ 49.47
Bob Stephenson  $ ∞

\underline{Person}  \underline{Description}  \underline{Approved}

George McNulty He’s a professor and claims to be a “humble servants”. I wonder how he can more than one servant.

sure

Doug Meade He’s just an ordinary servant but sometimes says more than he should.

not a chance

Person  Description  Approved

George McNulty He’s a professor and claims to be a “humble servants”. I wonder how he can more than one servant.

sure

Doug Meade He’s just an ordinary servant but sometimes says more than he should.

not a chance

Person  Description  Approved
George McNulty He’s a professor and claims to be a “humble servants”. I wonder how he can more than one servant. sure
Doug Meade  He’s just an ordinary servant but sometimes says more than he should.  not a chance
Person  Description  Approved

George McNulty He’s a professor and claims to be a “humble servants”. I wonder how he can more than one servant.

sure

Doug Meade He’s just an ordinary servant but sometimes says more than he should. not a chance